OpenStudy (anonymous):

(Help please) What is the simplified form of 6sqrX*6sqrX*6sqrX*6sqrX? (I know this is really easy I just forgot how to do it and need a refresher.)

OpenStudy (studygurl14):

\[6\sqrt{x}\times6\sqrt{x}\times6\sqrt{x}\times6\sqrt{x} = (6\sqrt{x})^4\] \[=6^4\sqrt{x^4}=1296x^2\]

OpenStudy (anonymous):

Can you explain it please because I forgot?

OpenStudy (anonymous):

Ok thanks I think I remember how to do it!

OpenStudy (studygurl14):

:)